Tuan deposited $1,500 in a saving account that pays an annual rate of 7% interest compounded semiannually. Adela deposited $1,500 in a savings account that pays 8% interest annually. How much more money will Adela have than Tuan at the end of 1 year?
Found 2 solutions by solver91311, lwsshak3:
Answer by solver91311(24713)



I already answered this one.

Tuan's account at the end of the year:



Adela's account at the end of the year:



Do the calculator work and compare the results.


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by lwsshak3(11628)


The best way to solve this problem is to use the compound interest formula,
A=P(1+R/n)^n,which states that the amount of interest (A) earned is equal to the principal(P) times one plus the interest per period raised to the number of compound periods(n)
For Tuan: P=1500(amt deposited),R=7% or .07, 2 compound periods
For Adela: P=1500(amt deposited),R=8% or .08, one period
Tuan earned = 1500(1+.07/2)^2=1606.84
Adela earned = 1500x(1+.08/1)^1=1620
Ans: Adela earned $13.16 more in interest than Tuan after one year.